# Existence of attracting periodic orbits in 3-dimensional strongly 2-cooperative systems

### v1

@@ IV Case Studies
We describe two case studies that demonstrate the application of Thm. 1 to well-known models in systems biology and chemistry.
IV-A 3-dimensional Goodwin oscillator model
@@
The Goodwin oscillator is a classical biochemical control circuit that regulates enzyme or protein synthesis by negative end-product feedback [14]. Consider the 3D version of the model,
@@
where $\alpha,\beta,\gamma>0$, and $m$ is a positive integer.
As noted by Gonze and Ruoff [13]: “The three-variable Goodwin model (adapted by Griffith) can be seen as a core model for a large class of biological systems, ranging from ultradian to circadian clocks.”
The state space of (19) is $\Omega:={\mathbb{R}}^{3}_{\geq 0}$. Furthermore, Griffith [15] showed that all trajectories are bounded, as any trajectory emanating from $\Omega$ eventually enters into the closed box
${\mathcal{B}}_{G}:=\{x\in{\mathbb{R}}^{3}_{\geq 0}\,|\,x_{1}\leq 1/\alpha,\;x_{2}\leq 1/(\alpha\beta),\;x_{3}\leq 1/(\alpha\beta\gamma)\}.$
@@
If $e$ is locally asymptotically stable, then we may expect that all solutions converge to $e$. Tyson [33] proved that the system (19) admits a periodic solution whenever $e$ is unstable, but provided no information on the stability of the periodic solution. Indeed, he states [33, p. 312]: “Notice that we are not proving that this closed orbit is a global attractor of the torus. Though we might expect this from the computer simulations, it would be much more difficult to prove than simple existence”.
We now considerably strengthen the result of [33] by showing that, whenever $e$ is unstable, the system admits at least one periodic orbit that is attracting. We also have consistency of oscillations with respect to the initial conditions: the solution emanating from any $x(0)\in{{\mathcal{B}}_{G}}\setminus\{e\}$ converges to a periodic orbit.
The Jacobian of (19) is
$J(x)=\begin{bmatrix}-\alpha&0&-\frac{mx_{3}^{m-1}}{(1+x_{3}^{m})^{2}}\\ 1&-\beta&0\\ 0&1&-\gamma\end{bmatrix},$
so the system is 2-cooperative on ${\mathcal{B}}_{G}$, as $J(x)$ has the sign pattern (2) . We now show that it is also strongly $2$-cooperative. Using (19), it is easy to verify that if $x(t)$ is a solution of (19) with $x(0)\in{\mathcal{B}}_{G}$ and $x_{3}(0)=0$, then there exists some $\delta>0$ such that $t\in(t_{0},t_{0}+\delta)\implies x_{3}(t)>0$. In particular, the set $\left\{t\geq 0\ |\ x_{3}(t)=0\right\}$ is at most countable, and this implies that $M(t)$ in (4), which is obtained from $J(x)$, is irreducible for almost all $t$.
The characteristic polynomial of $J(e)$ is
$\det(sI_{3}-J(e))$ $=s^{3}+(\alpha+\beta+\gamma)s^{2}+(\alpha\beta+\alpha\gamma+\beta\gamma)s$
$+\alpha\beta\gamma+(me_{3}^{m-1}/(1+e_{3}^{m})^{2}),$
so the free coefficient is positive. Thus, Thm. 1 applies to this case study and yields the following corollary.
@@
For the system (19) with $\alpha=0.5$, $\beta=0.4$, $\gamma=0.6$, and $m=10$, the box ${\mathcal{B}}_{G}$ becomes
${\mathcal{B}}_{G}=\{x\in{\mathbb{R}}^{3}_{\geq 0}\,|\,x_{1}\leq 2,\;x_{2}\leq 5,\;x_{3}\leq 25/3\},$
and the polynomial in (20) becomes
$Q(s)=0.12s^{11}+0.12s-1.$
The unique real and positive root of this polynomial is $e_{3}=1.1956$, so
$e=\begin{bmatrix}\beta\gamma e_{3}&\gamma e_{3}&e_{3}\end{bmatrix}^{\top}=\begin{bmatrix}0.2870&0.7174&1.1956\end{bmatrix}^{\top}.$
The characteristic polynomial is
$\det(sI_{3}-J(e))=s^{3}+1.5s^{2}+0.74s+1.1478,$
and applying the Routh stability criterion implies that $e$ is unstable. Indeed, the eigenvalues of $J(e)$ are $0.0062+\jmath 0.8711$, $0.0062-\jmath 0.8711$ and $-1.5125$, and correspond to the eigenvectors
@@
Fig. 2 depicts the solution $x(t)$ of (19) emanating from the initial condition $x(0)=\begin{bmatrix}0.1&0.1&0.1\end{bmatrix}^{\top}$. It can be seen that $x(t)$ converges to a periodic orbit.
@@
IV-B Field-Noyes model
The Field-Noyes ODE model for the Belousov-Zhabotinskii reaction, derived in [11], is given by
@@ IV-B Field-Noyes model
where $s,q,f,w$ are positive constants. The state variables represent concentrations in chemical reactions, so the state space is $\Omega:={\mathbb{R}}^{3}_{\geq 0}$.
We assume throughout that $q\ll 1$ (the numerical parameter value in [11] is $q=8.375\times 10^{-6}$, see also [16]). Then, the closed box
@@
with $y_{1}:=(1+q)^{-1}qf$ and $y_{2}:=(2q)^{-1}f$, is an invariant set of the dynamics (see [22]).
@@
$e_{1}$ $=\frac{1}{2q}\left(1-f-q+\left((1-f-q)^{2}+4q(1+f)\right)^{1/2}\right),$
$e_{2}$ $=\frac{e_{1}f}{1+e_{1}}=(1+f-qe_{1})/2,$ (22)
$e_{3}$ $=e_{1}.$
The Jacobian of (21) is
@@
and has the sign pattern
$J(x)=\begin{bmatrix}*&<0&0\\ <0&*&>0\\ >0&0&*\end{bmatrix},\text{ for all }x\in\operatorname{int}({{\mathcal{B}}}),$
implying that (21) is strongly 2-cooperative, up to a coordinate transformation [36]. A calculation shows that the free coefficient of the polynomial $\det(\lambda I_{3}-J(e))$ is
$a_{0}$ $=w(2qe_{1}^{2}+(2q+f-1)e_{1}+2e_{2}-f-1),$
and using (IV-B) gives
$a_{0}$ $=we_{1}(2qe_{1}+q+f-1)$
$=we_{1}\left((1-f-q)^{2}+4q(1+f)\right)^{1/2}>0.$
Thus, Thm. 1 applies to this case study and yields the following corollary.
@@
Consider the 3D Field-Noyes model (21) with equilibrium $e\in{\mathcal{B}}_{FN}$. Suppose that $J(e)$ is unstable. Then, for any initial condition $x(0)\in{\mathcal{B}}_{FN}\setminus\{e\}$, the solution of (21) converges to a periodic orbit.
@@
For the system (21) with $q=9.374\cdot 10^{-6}$, $f=1$, $s=0.3$, and $w=0.2934$, the box ${\mathcal{B}}_{FN}$ becomes
@@
where
$y_{1}=8.374\cdot 10^{-6},\ y_{2}=5.97\cdot 10^{4}.$
The corresponding equilibrium point is given by
$e=\begin{bmatrix}488.1780&0.9979&488.1780\end{bmatrix}^{\top}.$
The characteristic polynomial is
@@
so the free coefficient is positive, and since one of the coefficients is negative, $J(e)$ is unstable. Fig. 3 depicts the solution $x(t)$ of (21) emanating from the initial condition $x(0)=\begin{bmatrix}732.2670&9.9795&732.2670\end{bmatrix}^{\top}$. It can be seen that $x(t)$ converges to a periodic orbit.
Figure 3: A solution of the Field-Noyes system in Example 3 converges to a periodic orbit.
V Discussion
Strongly 2-cooperative systems enjoy an important asymptotic property: all bounded solutions that keep a positive distance from the set of equilibria converge to a periodic orbit. It is important to find cases where an attracting periodic orbit can be shown to exist. Here, we derived a simple sufficient condition that guarantees the existence of an attracting periodic orbit for 3D systems. The proof relies on the asymptotic and spectral properties of strongly 2-cooperative systems. Our result provides a unifying treatment of interesting case studies in systems biology and allows to derive stronger theoretical results than those available in the literature, while simplifying and unifying the proofs.
An $n$-dimensional system that is $(n-1)$-cooperative is , up to a coordinate transformation, a competitive system [35]. In particular, $2$-cooperative 3D systems are competitive systems. Indeed, let $D:=\mathop{\rm diag}\nolimits(1,-1,1)$ and $P:=\begin{bmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{bmatrix}$. Then for $n=3$ we have $\bar{A}_{2}=\begin{bmatrix}*&\geq 0&\leq 0\\ \geq 0&*&\geq 0\\ \leq 0&\geq 0&*\end{bmatrix}$, and
$DP\bar{A}_{2}(DP)^{-1}=\begin{bmatrix}*&\leq 0&\leq 0\\ \leq 0&*&\leq 0\\ \leq 0&\leq 0&*\end{bmatrix}.$
In other words, if $\dot{x}=Ax$ is a $2$-cooperative 3D system and $y(t):=DPx(t)$ then the $y$-system $\dot{y}=DPA(DP)^{-1}y$ is competitive. Thus, our results may also be interpreted as providing a sufficient condition guaranteeing that 3D competitive systems admit an attracting periodic orbit.
An interesting topic for further research is to find conditions that guarantee uniqueness of the periodic orbit. Then, our result would imply that the periodic orbit is a global attractor for solutions emanating from the invariant set ${\mathcal{H}}$. Another interesting research direction is to extend our result to systems with dimension $n>3$. Finally, our results may find application to the design of oscillators, which attracts considerable interest e.g. in synthetic biology [6, 23, 24].

### v2

@@ 4 Two Case Studies
We demonstrate the application of Theorem 1 to two well-known models in systems biology and chemistry.
4.1 3-dimensional Goodwin oscillator model
@@
The Goodwin oscillator is a biochemical circuit where enzyme or protein synthesis is regulated through the negative feedback of the end-product (Goodwin, 1965). Consider the 3D version of the model,
@@
where $\alpha,\beta,\gamma>0$, and $m$ is a positive integer. As noted by Gonza and Ruoff (2021): “The three-variable Goodwin model (adapted by Griffith) can be seen as a core model for a large class of biological systems, ranging from ultradian to circadian clocks.”
The state space of (19) is $\Omega:={\mathbb{R}}^{3}_{\geq 0}$, since the state variables represent concentrations of chemical species. Griffith (1968) showed that all trajectories are bounded, as any trajectory emanating from $\Omega$ eventually enters the closed box
${\mathcal{B}}_{G}:=\left\{x\in{\mathbb{R}}^{3}_{\geq 0}\,|\,x_{1}\leq\frac{1}{\alpha},\;x_{2}\leq\frac{1}{\alpha\beta},\;x_{3}\leq\frac{1}{\alpha\beta\gamma}\right\}.$
@@
If $e$ is locally asymptotically stable, then we expect that all solutions converge to $e$. Tyson (1975) proved that the system (19) admits a periodic solution whenever $e$ is unstable, but provided no information on convergence to a periodic solution. Indeed, he states (Tyson, 1975, p. 312): “Notice that we are not proving that this closed orbit is a global attractor of the torus. Though we might expect this from the computer simulations, it would be much more difficult to prove than simple existence”. We strengthen the result of Tyson (1975) by showing that, whenever $e$ is unstable, any solution emanating from $x(0)\in{{\mathcal{B}}_{G}}\setminus\{e\}$ converges to a periodic orbit.
System (19) is 2-cooperative on ${\mathcal{B}}_{G}$, because its Jacobian
$J(x)=\begin{bmatrix}-\alpha&0&-\frac{mx_{3}^{m-1}}{(1+x_{3}^{m})^{2}}\\ 1&-\beta&0\\ 0&1&-\gamma\end{bmatrix}$
has the sign pattern (2). The system is also strongly $2$-cooperative. In fact, if $x(t)$ is a solution of (19) with $x(0)\in{\mathcal{B}}_{G}$ and $x_{3}(0)=0$, then there exists some $\delta>0$ such that $t\in(0,\delta)\implies x_{3}(t)>0$. In particular, the set $\left\{t\geq 0\ |\ x_{3}(t)=0\right\}$ is at most countable, and this implies that $M(t)$ in (4), which is obtained from $J(x)$, is irreducible for almost all $t$.
Moreover, $\det(J(e))=-\alpha\beta\gamma+\frac{me_{3}^{m-1}}{(1+e_{3}^{m})^{2}}$ is negative. Thus, Theorem 1 yields the following corollary.
@@
For the system (19) with $\alpha=0.5$, $\beta=0.4$, $\gamma=0.6$, and $m=10$, the box ${\mathcal{B}}_{G}$ is ${\mathcal{B}}_{G}=\{x\in{\mathbb{R}}^{3}_{\geq 0}\,|\,x_{1}\leq 2,\;x_{2}\leq 5,\;x_{3}\leq 25/3\},$ and the polynomial in (20) is $Q(s)=0.12s^{11}+0.12s-1.$ The unique real and positive root of $Q$ is $e_{3}=1.1956$, so $e=\begin{bmatrix}\beta\gamma e_{3}&\gamma e_{3}&e_{3}\end{bmatrix}^{\top}=\begin{bmatrix}0.2870&0.7174&1.1956\end{bmatrix}^{\top}.$ The characteristic polynomial is $\det(sI_{3}-J(e))=s^{3}+1.5s^{2}+0.74s+1.1478,$ and applying the Routh stability criterion implies that $e$ is unstable. Indeed, the eigenvalues of $J(e)$ are $0.0062+\jmath 0.8711$, $0.0062-\jmath 0.8711$ and $-1.5125$, with corresponding eigenvectors
@@
Fig. 2 depicts the solution $x(t)$ of (19) emanating from the initial condition $x(0)=\begin{bmatrix}0.1&0.1&0.1\end{bmatrix}^{\top}$, which converges to a periodic orbit.
@@
4.2 Field-Noyes model
The ODE model for the Belousov-Zhabotinskii reaction derived by Field and Noyes (1974) is given by
@@ 4.2 Field-Noyes model
where $s,q,f,w$ are positive constants. The state variables represent concentrations of species involved in chemical reactions, so the state space is $\Omega:={\mathbb{R}}^{3}_{\geq 0}$.
@@
[…]
with $y_{1}:=(1+q)^{-1}qf$ and $y_{2}:=(2q)^{-1}f$, is an invariant set for the dynamics (see Murray (1974)).
@@
$\begin{cases}e_{1}=\frac{1-f-q+\sqrt{(1-f-q)^{2}+4q(1+f)}}{2q},\\ e_{2}=\frac{e_{1}f}{1+e_{1}}=\frac{1+f-qe_{1}}{2},\\ e_{3}=e_{1}.\end{cases}$ (22)
The Jacobian of (21),
@@
has the sign pattern $\begin{bmatrix}*&<0&0\\ <0&*&>0\\ >0&0&*\end{bmatrix}$ for all $x\in\operatorname{int}({{\mathcal{B}}})$, implying that (21) is strongly 2-cooperative, up to a coordinate transformation (Weiss and Margaliot, 2021b). Computing the determinant of the Jacobian and substituting the equilibrium value in (22) gives
$\det(J(e))$ $=-w(2qe_{1}^{2}+(2q+f-1)e_{1}+2e_{2}-f-1)$
$=-we_{1}(2qe_{1}+q+f-1)$
$=-we_{1}\sqrt{(1-f-q)^{2}+4q(1+f)}<0,$
and applying Theorem 1 yields the following result.
@@
Consider the Field-Noyes model (21) with equilibrium $e\in{\mathcal{B}}_{FN}$. Suppose that $J(e)$ is unstable. Then, for any initial condition $x(0)\in{\mathcal{B}}_{FN}\setminus\{e\}$, the solution of (21) converges to a periodic orbit.
@@
For the system (21) with $q=9.374\cdot 10^{-6}$, $f=1$, $s=0.3$, and $w=0.2934$, the box ${\mathcal{B}}_{FN}$ is
@@
where $y_{1}=8.374\cdot 10^{-6}$ and $y_{2}=5.97\cdot 10^{4}$. The corresponding equilibrium is $e=\begin{bmatrix}488.1780&0.9979&488.1780\end{bmatrix}^{\top}.$ The characteristic polynomial is
@@
so $\det(J(e))=-1.1722$, and since one of the coefficients is negative, $J(e)$ is unstable. Fig. 3 depicts the solution $x(t)$ of (21) emanating from the initial condition $x(0)=\begin{bmatrix}732.2670&9.9795&732.2670\end{bmatrix}^{\top}$, and shows that $x(t)$ converges to a periodic orbit.
Figure 3: Solution of the Field-Noyes system in Example 2.
5 Discussion
Strongly 2-cooperative systems satisfy an important asymptotic property: all bounded solutions that keep a positive distance from the equilibrium set converge to a periodic orbit. Here, we derived a simple sufficient condition for 3D systems that guarantees the existence of an invariant set $\mathcal{G}$ that includes no equilibrium points, such that all solutions emanating from $\mathcal{G}$ converge to a non-trivial periodic orbit, and we provide an explicit characterization of the set $\mathcal{G}$. The proof relies on the asymptotic and spectral properties of strongly 2-cooperative systems.
An $n$-dimensional system that is $(n-1)$-cooperative is, up to a coordinate transformation, a competitive system (Weiss and Margaliot, 2021a). In particular, $2$-cooperative 3D systems are competitive systems. Thus, our results may also be interpreted in the framework of 3D competitive systems (Smith, 1995). However, we believe that our proposed approach can be generalized to strongly $2$-cooperative systems of general dimension $n$, which are not competitive systems. This topic is currently under study.
Another interesting research direction is applications to the design of oscillators, which attracts considerable interest e.g. in synthetic biology (Blanchini et al., 2014; Novák and Tyson, 2008; Panghalia and Singh, 2020).
Acknowledgement
We thank the anonymous reviewers for their helpful comments.